OpenStudy (anonymous):

Which value is a solution for the equation tan(x/2)=1? A. π/2 B. 3π/2 C. 3π/4 D. π/4

OpenStudy (xapproachesinfinity):

tanx=1, x=phi/4+kpi

OpenStudy (xapproachesinfinity):

use this fact to solve tan(x/2)=1

OpenStudy (anonymous):

Wait, do you mean x=pi/4+kpi or do you actually mean the "phi"?

OpenStudy (xapproachesinfinity):

no what i mean is the values that can make tangent 1 are pi/4+kpi this a general solution. you need to derive you solution from this

OpenStudy (xapproachesinfinity):

your x here is x/2! do you get it

OpenStudy (anonymous):

Yes, I get it :o Give me a moment to work it out

OpenStudy (xapproachesinfinity):

Okay!

OpenStudy (anonymous):

Hm, I'm actually getting sort of confused. :< Would I plug in "pi/4+kpi" into the "x" in tan(x/2)=1?

OpenStudy (xapproachesinfinity):

No, tan(x/2)=1 means that x/2=pi/4+kpi solve for x

OpenStudy (anonymous):

Ahhh, what would be done with "k", then?

OpenStudy (xapproachesinfinity):

i'm just showing you that if tan(x)=1 that means x must be x=pi/4+kpi x being an angle! so the same reasoning there some angles such that tan(x/2)=1 we know that those angles much be pi/4+kpi (k is the period) so x/2=pi/4+kpi

OpenStudy (xapproachesinfinity):

just leave it

OpenStudy (xapproachesinfinity):

actually you question should state something about the from ?<x<? since they give you the choices what we are doing is the general solution!!

OpenStudy (xapproachesinfinity):

@nincompoop what do think lol

OpenStudy (anonymous):

Hm, I'm still getting rather lost x) Would it be closer to pi/4 as a solution, then?

OpenStudy (nincompoop):

oh I was not reading, I accidentally clicked it and I just didn't bother

OpenStudy (xapproachesinfinity):

haha well x/2=pi/4+kpi how do you solve for x here? do you know how

OpenStudy (xapproachesinfinity):

@nincompoop oh, alright! xD

OpenStudy (xapproachesinfinity):

did your question give interval for x? like 0<x<pi/2 or something

OpenStudy (nincompoop):

you will have to think tangent in terms of sin/cosine

OpenStudy (nincompoop):

oh wait you have options

OpenStudy (xapproachesinfinity):

well even with that, those options are not accurate if there is not interval?

OpenStudy (xapproachesinfinity):

that equation will simply be solved in general terms?

OpenStudy (xapproachesinfinity):

tangent is positive in two quadrants

OpenStudy (nincompoop):

yeah you can use tan = sin/cos concept

OpenStudy (nincompoop):

sin = y cos = x x=y = 1 your slope is at a 45 degree angle which is 180/4 or pi/4

OpenStudy (xapproachesinfinity):

well that's pi/4+kpi no?

OpenStudy (xapproachesinfinity):

depends on the period

OpenStudy (nincompoop):

yeah, but that is your basis

OpenStudy (xapproachesinfinity):

he hasn't given any other info to use to solve for a specific interval that's why im saying the solution is in general temrs?

OpenStudy (xapproachesinfinity):

wondered why there are options like that xD

OpenStudy (nincompoop):

x = pi/2 = 90 90/2 = 45

OpenStudy (nincompoop):

tan 45 = 1

OpenStudy (xapproachesinfinity):

well tan225=1 too? haha

OpenStudy (xapproachesinfinity):

that's 5pi/4

OpenStudy (nincompoop):

yes, but can you get tan 225 with any of the options? I doubt

OpenStudy (xapproachesinfinity):

anyway for this question the solution is pi/2. but the question is not accurate as i said perhaps there is a reason

OpenStudy (nincompoop):

it is accurate, but slightly tricky

OpenStudy (xapproachesinfinity):

But let's talk in general! this equation as it is has more than one solution without considering options

OpenStudy (xapproachesinfinity):

time to sleep haha not for trig xD